OpenStudy (tmagloire1):

An ice block is melting at the rate of 1.5 inches per hour. How fast is the surface area of the ice cube changing at the instant the ice block has a side length of 2 inches? −12 in2/hour −18 in2/hour −36 in2/hour −24 in2/hour

OpenStudy (welshfella):

if side lenght = s Area A = 6s^2 ds/dt = 1.5 dA/ds = 12x = 24 at s = 2 Rate of change of surface area at s = 2 = dA/dt is dA/dt = ds/dt *dA/ds

OpenStudy (tmagloire1):

Okay but how do I find dA/dt @welshfella

OpenStudy (welshfella):

check my post i found dA/ds in that

OpenStudy (welshfella):

dA/dt = ds/dt * dA/ds = 1.5 * 24

OpenStudy (tmagloire1):

Ohhh so it would be 1.5(24)

OpenStudy (welshfella):

yea

OpenStudy (tmagloire1):

But - because it's falling?

OpenStudy (welshfella):

oh right my mistake its -36

OpenStudy (tmagloire1):

Okay I understand thank you for your help!

OpenStudy (welshfella):

yw